# Mechanism Shift During Post-training from Autoregressive to Masked Diffusion Language Models

### v2

@@ 4.2 Semantic Reorganization of Task-Critical Components and Early Layers
To understand how circuit-level differences translate into concrete computational strategies, we analyze model behavior at two complementary granularities. First, we examine task-critical components—attention heads and MLPs that receive high EAP-IG attribution—using component-wise logit lens analysis. This allows us to characterize how semantic roles are assigned to components that directly influence task outputs. Second, we analyze early-layer neuron activations to understand how these semantic roles are implemented internally, particularly in regions where MDMs emphasize computation but component-level semantics appear diffuse.
We begin by analyzing components that directly contribute to task performance, as identified by high EAP-IG attribution. By applying a component-wise logit lens to these components, we probe their semantic alignment with output tokens, allowing us to assess whether task-relevant information is concentrated within a small number of specialized components or distributed across many.
Causal reasoning task (IOI)
For the IOI task, ARMs exhibit sharply localized semantic specialization. In both Qwen and LLaMA, a small number of attention heads are strongly aligned with person-name tokens (e.g., Dan, Jerry), producing large-magnitude logits that dominate prediction. This pattern indicates that IOI is largely solved by a limited set of highly specialized components that act as deterministic, pointer-like mechanisms.
DiffuLLaMA largely preserves this behavior. Several high-attribution components remain strongly aligned with person-name tokens and exhibit logit profiles comparable to those of the underlying ARM. This suggests substantial inheritance of autoregressive circuitry, consistent with the observation that IOI relies primarily on local causal dependencies that are already well supported by the autoregressive objective.
Dream, however, departs from this pattern. Its high-attribution components do not exhibit clear person-name specificity. Instead, the strongest logit alignments correspond to non-person proper nouns or broader semantic tokens (e.g., Browser, Kremlin). This indicates that, despite similar circuit topology at a coarse level, diffusion post-training redistributes semantic roles across components, weakening the dominance of individual name-specific heads.
Global reasoning task (Countdown).
A contrasting pattern emerges in the Countdown task. ARMs rely on components with strong numerical selectivity, often sharply aligned with specific digits or operators. These components produce high-confidence logits concentrated on a narrow subset of tokens, indicating that global planning is approximated through sequential, component-centric heuristics.
@@ 4.2.1 Task-Critical Components: Component-wise Logit Lens Analysis
In Dream, this sharp specialization collapses. Instead of a single dominant numerical component, multiple components exhibit moderate logit responses distributed across numerical tokens, with no individual component exerting decisive control over prediction. DiffuLLaMA again occupies an intermediate regime, retaining numerical associations but with reduced magnitude and increased dispersion.
@@
Taken together, these results indicate that diffusion post-training induces a semantic reorganization at the component level. Whereas ARMs resolve tasks through a small number of highly specialized components, MDMs—particularly on global reasoning tasks—redistribute semantic responsibility across multiple components, yielding a more ensemble-like computation.
@@ 4.2.2 Early-layer Representation: Neuron-level Analysis
While component-wise logit lens analysis captures semantic specialization among task-critical components, it does not fully explain the behavior of early-layer regions that are emphasized by masked diffusion circuits. To characterize how task-relevant information is implemented in these layers, we analyze neuron activations in the lowest transformer layers across tasks and model classes.
In ARMs, neuron activation in early layers is highly task-dependent. In the IOI task, a large number of neurons in early layers are strongly activated by descriptive adjectives and modifier-related tokens, suggesting that early layers encode fine-grained syntactic and descriptive features prior to resolving entity identity. In the Countdown task, this pattern shifts markedly toward neurons associated with numerical or technical content. Despite this task-dependent reallocation, ARMs consistently exhibit strong concentration within specific semantic categories.
MDMs display a qualitatively different pattern. Across both IOI and Countdown, the total number of active neurons in early layers remains relatively stable. Moreover, the activated neurons tend to correspond to broad, genre-level cues—such as general numerical or technical content—rather than sharply defined task-specific categories. This indicates that MDMs rely less on early-layer specialization and instead maintain a more uniform, task-agnostic representational regime.
This contrast persists in the early layer. Across depth, ARMs repeatedly exhibit sharper category-level specialization, while MDMs maintain flatter activation profiles with fewer neurons strongly associated with any single semantic category.
@@ 4.2.3 Connecting Component-level and Neuron-level Perspectives
Component-wise logit lens and neuron-level visualization serve complementary interpretability roles. Logit lens probes semantic alignment at the level of components, revealing which tokens individual attention heads or MLPs are linearly aligned with and how they directly influence model outputs. Neuron-level analysis, by contrast, characterizes how semantic information is internally implemented, independent of direct alignment with the output vocabulary.
Together, these analyses reveal a consistent mechanistic pattern that is further supported by quantitative differences in explanation structure. ARMs exhibit a smaller number of unique explanatory components (266) but substantially higher variance in explanations (), indicating that model behavior is dominated by a limited set of highly influential and specialized components. In contrast, MDMs rely on a larger pool of explanatory components (426) with markedly lower variance (), suggesting that explanatory responsibility is distributed more evenly across components.
This quantitative shift aligns with the qualitative patterns observed in both analyses. ARMs employ a component-centric strategy in which task resolution is dominated by a small number of sharply specialized attention heads, or MLPs, supported by task-specific early-layer neurons. MDMs, by contrast, replace these sharp semantic roles with distributed responsibility, activating broader sets of components whose individual contributions are weaker but collectively stable. At the neuron level, this manifests as early-layer representations that exhibit reduced task-specific selectivity and more uniform activation profiles across tasks.
Taken together, these findings provide a coherent mechanistic account of diffusion post-training. MDMs trade component-level specialization for distributed semantic coverage, supported by early layers that provide broad, task-invariant representational scaffolding. This reorganization explains both the front-loaded computation observed in circuit analyses and the absence of dominant, easily interpretable components in logit-lens probes, highlighting a fundamental shift from localized causal pathways to globally integrated computation.
@@ 5 Conclusion
In this paper, we investigated how post-training an autoregressive language model (ARM) with a masked diffusion objective reshapes its internal computational mechanisms. By combining circuit analysis, component-wise logit-lens probing, and neuron-level explanaion, we provided a mechanistic comparison between ARMs and their post-trained masked diffusion model (MDM) counterparts across tasks with distinct structural demands.
@@
Our analysis focuses on a limited set of tasks—IOI and Countdown—which are chosen to isolate causal dependency and global planning behaviors, respectively. While these tasks are well-established benchmarks in mechanistic interpretability, the circuits identified in this work should be interpreted as task-specific mechanisms rather than universal properties of autoregressive or masked diffusion language models. Different linguistic or reasoning tasks may recruit alternative circuitry or induce distinct patterns of post-training adaptation. In addition, our circuit discovery pipeline relies on sparsity-inducing methods such as EAP-IG and Top-K divergence selection to isolate the most behaviorally salient components. Although this design choice improves interpretability by filtering out noisy or redundant pathways, it implies that not all attention heads and MLP components are exhaustively analyzed, and secondary or auxiliary mechanisms that contribute weakly to task performance may be omitted. Finally, our approach involves an inherent trade-off between interpretability resolution and computational scalability: component-wise logit lens analysis and neuron-level visualization across diffusion steps are computationally expensive, particularly for large-scale models. As a result, we prioritize in-depth analysis of selected components over exhaustive full-model sweeps, and extending the analysis to broader component sets or additional configurations would require substantially greater computational resources.
Appendix A Dataset Details
@@
For the countdown task, we evaluate both LLaMA-series and Qwen-series models using 500 evaluation examples per model. For LLaMA-series models, we use 13 diffusion steps, which correspond to 13 tokens (diffusion step = token). For Qwen-series models, we use 12 diffusion steps, corresponding to 12 tokens. For the IOI task, we evaluate 500 examples in a single-step setting with 1 diffusion step and 1 token.
Appendix B Experimental Details and Results
@@
Our experiments use HuggingFace Transformers for model loading and inference. Circuit discovery is implemented using Edge Attribution Patching with Integrated Gradients (EAP-IG), where we select the top 1,000 edges(corresponding to a faithfulness score of 0.6) ranked by attribution scores as a trade-off between sparsity and attribution coverage. Empirically, smaller thresholds lead to unstable circuit structures, while larger thresholds rapidly approach the full graph without improving interpretability. The choice of 1,000 edges lies in a stable regime where circuit topology and component rankings remain consistent.
Attribution scores are then aggregated at the component level to identify the Top-K components. We set , corresponding to the smallest value for which the selected components form a well-connected circuit, ensuring that the resulting subgraph remains interpretable while preserving end-to-end connectivity. Logit lens projections follow the standard unembedding-based formulation. All other parameters follow default settings from the respective libraries.
@@
All pretrained models and tools used in this work are publicly released research artifacts. We use them solely for research and analysis purposes, in accordance with their respective licenses, and do not redistribute any models or derived data.
@@
For analyses involving step-wise circuit extraction, we compute circuits at each diffusion step and then aggregate attribution scores across steps. We find that the set of participating components (attention heads and MLP layers) remains largely stable over diffusion time, with more than 90% overlap in selected components between steps. Accordingly, the step-wise circuit visualizations shown in Figures 6 and 5 represent averaged structures, while step-wise variation is primarily reflected in changes to the attributed edges rather than in the identity of the components themselves. This suggests that masked diffusion primarily refines information routing among a largely fixed component set, rather than progressively recruiting new components.

### v3

@@ 4.2 Semantic Reorganization of Task-Critical Components and Early Layers
| Category | Dream | DiffuLLaMA | Qwen | LLaMA |
| PERSON | 1,300 | 337 | 298 | 528 |
| OTHER | 8,348 | 3,251 | 2,965 | 14,606 |
| ORG | 62 | 30 | 81 | 418 |
| LOC | 55 | 33 | 31 | 343 |
| DIGIT | 32 | – | 2 | 55 |
| OPERATOR | 20 | 2 | 210 | 124 |
| MISC | 3 | 2 | 2 | – |
| Total | 9,820 | 3,933 | 3,775 | 15,619 |
| NameFrac@10 | 13.2% | 8.6% | 7.9% | 3.4% |
To understand how circuit-level differences translate into concrete computational strategies, we analyze model behavior at two complementary granularities. First, we examine task-critical components—attention heads and MLPs that receive high EAP-IG attribution—using component-wise logit lens analysis. This allows us to characterize how semantic roles are assigned to components that directly influence task outputs. Second, we analyze early-layer neuron activations to understand how these semantic roles are implemented internally in regions where MDMs emphasize computation but component-level semantics appear diffuse.
@@ 4.2.1 Task-Critical Components: Component-wise Logit Lens Analysis
We analyze task-critical components identified by high EAP-IG attribution using a component-wise logit lens. This allows us to test whether task-relevant information is concentrated in a few specialized components or distributed across many. We report probes of semantic alignment and dominance in the main text, and defer representative top-token examples to Table 3, LABEL:tab:logit_lens.
Causal reasoning task (IOI):
For IOI, ARMs exhibit sharply localized semantic specialization. This pattern appears in the way person-name tokens are concentrated among the aligned outputs of task-critical components. As shown in Table 2, MDMs can place person-name tokens relatively often among their top aligned candidates—Dream achieves a higher (13.2%) than Qwen (7.9%) and LLaMA (3.4%), while DiffuLLaMA reaches 8.6%—but this does not imply sharper specialization. Instead, ARMs exhibit substantially stronger selective amplification once a relevant name token appears. Qwen and LLaMA achieve median LME values of 0.90 and 0.89, respectively, compared with 0.31 for Dream and 0.05 for DiffuLLaMA (Appendix Table 9). ARMs also show larger logit gaps, while DiffuLLaMA exhibits higher entropy, consistent with a more distributed alignment pattern. In short, raw name frequency does not imply sharper specialization: ARMs remain dominated by a few decisive name-specific components, whereas MDMs spread name-related evidence across a broader set of components.
This qualitative difference is also visible in representative component-level examples (Table 3). In Qwen and LLaMA, several high-attribution attention heads align strongly with person-name tokens, consistent with highly specialized, pointer-like behavior. DiffuLLaMA preserves this pattern, indicating substantial inheritance of autoregressive circuitry. Dream, however, departs from it: its high-attribution components are less clearly name-specific, and their strongest alignments often correspond to non-person proper nouns or broader semantic tokens. Thus, even when circuit topology is similar, diffusion post-training can redistribute semantic roles across components and weaken the dominance of individual name-specific heads.
Global reasoning task (Countdown):
@@
A contrasting pattern emerges in Countdown. ARMs rely on components with strong numerical selectivity, sharply aligned with specific digits or operators, suggesting that global planning is approximated through sequential, component-centric heuristics (Table 3). In Dream, this sharp specialization collapses: instead of a single dominant numerical component, multiple components exhibit moderate responses across numerical tokens with no individual component exerting decisive control. DiffuLLaMA occupies an intermediate regime, retaining numerical associations but with reduced magnitude and increased dispersion. This redistribution is also architectural—Dream’s CoG on Countdown drops from 16.5 (Qwen) to 4.8 (Table 1), indicating that diffusion post-training front-loads task-relevant computation toward earlier layers.
@@
Taken together, these results indicate that diffusion post-training induces a semantic reorganization at the component level. Whereas ARMs resolve tasks through a small number of highly specialized components, MDMs redistribute semantic responsibility across multiple components, yielding a more ensemble-like computation—accompanied in Dream by a pronounced shift toward earlier layers.
@@ 4.2.2 Early-layer Representation: Neuron-level Analysis
While component-wise logit lens analysis captures semantic specialization among task-critical components, it does not fully explain the behavior of early-layer regions emphasized by MDM circuits. To characterize how task-relevant information is implemented there, we analyze neuron activations in the lowest transformer layers.
This analysis is especially important for Countdown, where the CoG shift in Table 1 indicates that post-training moves substantial computation into the bottom of the network. In ARMs, early-layer neuron activation is highly task-dependent. In IOI, many early-layer neurons are strongly activated by descriptive adjectives and modifier-related tokens (e.g., 2,821 neurons in LLaMA; Appendix Table 5), suggesting these layers encode syntactic features before resolving entity identity. In Countdown, this pattern shifts toward neurons associated with numerical or technical content (over 1,000 neurons). Despite this task-dependent reallocation, ARMs consistently exhibit strong concentration within specific semantic categories.
MDMs display a qualitatively different pattern. Across both tasks, the number of active neurons in early layers remains relatively stable (typically 50–200 active neurons per top category; Appendix Table 5), and these neurons tend to correspond to broad, genre-level cues rather than sharply defined task-specific categories. This suggests MDMs rely less on category-specific early-layer specialization and instead maintain a more uniform, task-agnostic representational regime. Combined with the CoG results, this pattern implies that MDM early layers absorb more computation without relying on sharply specialized neurons.
@@ 4.2.3 Connecting Component-level and Neuron-level Perspectives
Taken together, the component-level and neuron-level analyses reveal a consistent mechanistic pattern. Component-wise logit lens shows how task-relevant outputs are distributed across influential components, whereas neuron-level analysis shows how these representations are implemented in the early layers. More specifically, logit lens probes semantic alignment at the component level, revealing which tokens individual attention heads or MLPs are linearly aligned with and how they influence model outputs. Neuron-level analysis, by contrast, characterizes how semantic information is internally implemented, independent of direct alignment with the output vocabulary.
In ARMs, explanatory mass is concentrated in a few highly influential components, whereas in MDMs it is distributed across a broader pool of weaker but collectively supportive components. ARMs exhibit a smaller number of unique explanatory components (266) but a higher standard deviation in explanations (), indicating that model behavior is dominated by a limited set of highly influential and specialized components. In contrast, MDMs rely on a larger pool of explanatory components (426) with lower standard deviation (), suggesting that explanatory responsibility is distributed more evenly across components.
ARMs employ a component-centric strategy in which task resolution is dominated by a small number of sharply specialized attention heads or MLPs, supported by task-specific early-layer neurons. MDMs, by contrast, replace these sharp semantic roles with distributed responsibility, activating broader sets of components whose individual contributions are weaker but collectively stable. At the neuron level, this appears as early-layer representations with reduced task-specific selectivity and more uniform activation profiles across tasks.
Taken together, these findings provide a mechanistic account of diffusion post-training. MDMs trade component-level specialization for distributed semantic coverage, supported by early layers that provide broad, task-invariant representational scaffolding. This reorganization explains both the front-loaded computation observed in circuit analyses and the reduced dominance of interpretable components in logit-lens probes, highlighting a fundamental shift from localized causal pathways to more globally integrated computation.
@@ 5 Conclusion
In this paper, we investigated where and how masked diffusion post-training reshapes the internal mechanisms of ARMs. To identify where, we use circuit analysis and layer-wise CoG to trace changes in circuit composition and depth-wise computation. To understand how, we use component-wise logit-lens probing and neuron-level analysis to examine shifts in semantic role assignment and representation patterns. Our results show that masked diffusion post-training induces a task-dependent mechanism shift rather than merely reusing autoregressive computation. In IOI, MDMs largely preserve the core autoregressive circuit while replacing sharp specialization with distributed semantic responsibility. In Countdown, MDMs show broader circuit rewiring, a stronger shift toward earlier layers, and the same transition to distributed component roles. Overall, masked diffusion post-training preserves autoregressive mechanisms for local causal reasoning, while reorganizing them for more globally integrated, non-sequential reasoning.
@@
Our study is confined to two benchmark tasks, IOI and Countdown, selected to capture two core yet contrasting modes of computation central to our inquiry: locally causal dependency tracking and global sequence-level planning. This pairing enables a controlled comparison of whether post-trained masked diffusion models preserve autoregressive mechanisms or develop distinct computational circuitry under differing structural demands. Nevertheless, the circuits identified here should be understood as task-specific findings rather than universal signatures of autoregressive or masked diffusion language models. Establishing broader generality will require evaluation across additional tasks spanning diverse compositional, linguistic, and reasoning structures. Furthermore, while our attribution-guided pipeline enables targeted mechanistic analysis by isolating the most behaviorally salient edges and components, it does not constitute an exhaustive characterization of all computations underlying model behavior. A more complete account would necessitate wider circuit sweeps across a broader range of tasks, models, and training conditions — an endeavor that remains computationally prohibitive at scale.
Appendix A Task Details
@@
For Countdown, we evaluate both LLaMA-series and Qwen-series models on 500 examples per model. LLaMA-series models use 13 diffusion steps (13 tokens), while Qwen-series models use 12 diffusion steps (12 tokens). For IOI, we evaluate 500 examples in a single-step setting with 1 diffusion step and 1 token. Since IOI in our setup requires generating only a single target token, increasing the number of diffusion steps would still correspond to repeatedly refining the same one-token prediction. As a result, the underlying circuit is not expected to change qualitatively across multiple steps, making the single-step setting the most direct and interpretable choice.
Dream-based conditional-entropy criterion
Following Section 3, we operationalize the local–global reasoning criterion using Dream, which yields the most stable and interpretable conditional-entropy estimates among the MDMs we study. For a target token , we compare prefix-only entropy, , with full-context entropy excluding the target token, . A larger positive gap indicates stronger global reasoning, while a smaller or negative gap indicates a more local or prefix-dominant regime.
Entropy-gap results
Table 4 reports Dream-based conditional entropy for the two tasks. Countdown shows a positive entropy gap, indicating that full-sequence context substantially reduces uncertainty relative to prefix-only conditioning. In contrast, IOI shows a negative gap, consistent with a more local or prefix-dominant regime.
| Task | Prefix-only | Full-context | Gap |
| Countdown | 1.2756 [1.2517, 1.2984] | 0.8552 [0.7879, 0.9202] | +0.4204 [0.3456, 0.4911] |
| IOI | 3.4383 [3.3125, 3.5644] | 4.1884 [4.0699, 4.3050] | -0.7500 [-0.9237, -0.5783] |
Appendix B Experimental Details and Additional Results
@@
We use HuggingFace Transformers for model loading and inference. Circuit discovery is implemented with Edge Attribution Patching with Integrated Gradients (EAP-IG). We select the top 1,000 edges, corresponding to a faithfulness score of 0.6, as a trade-off between sparsity, attribution coverage, and graph connectivity. In practice, reducing the threshold below 1,000 leads to fragmented circuits: we observe that at 800 edges, the discovered subgraph no longer preserves end-to-end connectivity. By contrast, increasing the threshold yields graphs that more closely approach the full computation graph without providing additional interpretive benefit. The choice of 1,000 edges therefore lies in a stable regime where circuit topology, connectivity, and component rankings remain consistent.
Attribution scores are then aggregated at the component level to identify the Top-K components. We set , the smallest value for which the selected components form a well-connected and interpretable subgraph while preserving end-to-end connectivity. Importantly, these top 100 components account for more than 65% of the total attribution mass, indicating that a relatively small subset of components captures most of the task-relevant computation. Logit-lens projections follow the standard unembedding-based formulation. All other parameters use default library settings.
@@
All pretrained models and tools used in this work are publicly released research artifacts. We use them solely for research and analysis in accordance with their respective licenses, and do not redistribute any models or derived data.
@@
For step-wise circuit extraction, we compute circuits at each diffusion step and aggregate attribution scores across steps. The set of participating components (attention heads and MLP layers) remains largely stable over diffusion time, with more than 70% overlap in selected components between steps. Accordingly, the visualizations in Figures 7 and 6 represent averaged structures; step-wise variation is reflected mainly in the attributed edges rather than in component identity. This suggests that masked diffusion primarily refines information routing within a largely fixed component set, rather than progressively recruiting new components.
| Task | Model | Layer | Explanation | Count |
| IOI | LLaMA | 0–4 | descriptive adjectives | 2821 |
| expressions of uncertainty or doubt | 104 | |||
| synonyms for happiness or joy | 50 | |||
| DiffuLLaMA | 0–4 | synonyms for happiness or joy | 193 | |
| expressions of uncertainty or doubt | 80 | |||
| technical jargon or specialized terminology | 57 | |||
| Countdown | LLaMA | 0–4 | numerical values or quantities | 1012 |
| terms associated with technology and innovation | 226 | |||
| technical terms related to technology | 86 | |||
| DiffuLLaMA | 0–4 | synonyms for happiness or joy | 207 | |
| expressions of uncertainty or doubt | 114 | |||
| terms associated with technology and innovation | 72 |
@@
Appendix C Quantitative Mechanism Metrics & Results
To formalize specialization and dominance in our component-level logit lens analysis, and to provide objective support for our qualitative interpretations, we introduce several quantitative probes. Let denote the logit assigned by component to token , and let denote the set of top- tokens ranked by logit.
C.1 Semantic Alignment and Dominance Probes
Name Alignment Frequency (NameFrac@K):
To quantitatively assess whether semantic alignment extends beyond the top-1 token, we measure the fraction of person-name tokens among the top- aligned tokens for each component:
where denotes the set of person-name tokens, identified using a pretrained BERT-based Named Entity Recognition (NER) model (filtered for the PERSON label).
Selective Amplification (Logit Gap):
To quantify whether a component strongly favors a single token, we compute the logit gap between the highest and second-highest logits. Because softmax probabilities depend exponentially on logit differences, larger logit gaps indicate stronger selective amplification:
where and represent the maximum and second-maximum logits, respectively.
Log-Mean-Exp Dominance Gap (LME):
Because frequency alone does not measure predictive dominance, we quantify whether name tokens exert stronger influence when they appear by computing a log-mean-exp dominance gap between name and non-name tokens within the top- set:
where the LME for a subset of tokens is defined as:
This metric controls for group size and directly measures selective amplification.
Distributional Sharpness (Entropy):
To formalize whether semantic alignment is concentrated or distributed, we compute the entropy over the normalized top- logits:
Lower entropy indicates sharper specialization (typical of autoregressive models), while higher entropy indicates more distributed alignment (typical of masked diffusion models).
C.2 Architectural Depth Metric
Center of Gravity (CoG):
To provide an objective, scalar measure of the “Mechanism Shift” independent of graph layout and visual density, we calculate the Center of Gravity (CoG). The CoG represents the attribution-weighted average layer index of the discovered circuit, identifying where the core computations are localized:
where is the layer index and is the sum of EAP-IG attribution scores for all components in layer .
The tables below provide the empirical results derived from the metrics defined above. We observe that autoregressive models exhibit larger logit gaps (e.g., Qwen: 2.12; LLaMA: 0.70) than their masked diffusion counterparts (DiffuLLaMA: 0.49), indicating a stronger concentration of probability mass on a single dominant token. Furthermore, DiffuLLaMA exhibits higher entropy (2.01) than autoregressive models (LLaMA: 1.81; Qwen: 0.97), indicating more distributed semantic alignment. Together, these metrics demonstrate the differences in specialization, selective amplification, and architectural depth between ARMs and MDMs.
| Model | Name Token Proportion (Top-1) | Median LME | Logit Gap | Entropy |
| Dream | 19.7% | 0.31 | – | – |
| DiffuLLaMA | 25.2% | 0.05 | 0.49 | 2.01 |
| Qwen | 6.0% | 0.90 | 2.12 | 0.97 |
| LLaMA | 13.4% | 0.89 | 0.70 | 1.81 |
| Task | Model | CoG (Layer Index) | Relative CoG (0–1) |
| IOI | Qwen2.5-7B | 17.528 | 0.548 |
| Dream-Base-7B | 20.432 | 0.638 | |
| Countdown | Qwen2.5-7B | 16.582 | 0.518 |
| Dream-Base-7B | 4.856 | 0.152 |